OpenStudy (anonymous):

if f(x) is a linear func. such that f(2)=5 & f(4)=13, f(x)=

OpenStudy (gorv):

linear function mean ax+b

OpenStudy (gorv):

f(x)=ax+b

OpenStudy (gorv):

to find a and b u r given f(2) and f(4)

OpenStudy (gorv):

put x=2

OpenStudy (gorv):

f(2)=5=a*2+b

OpenStudy (gorv):

5=2a+b

OpenStudy (gorv):

f(4)=13=a*4+b

OpenStudy (gorv):

13=4a+b

OpenStudy (gorv):

can u solve these two equation